### v1

# xRAG: Extreme Context Compression for Retrieval-augmented Generation with One Token

Abstract
This paper introduces xRAG, an innovative context compression method tailored for retrieval-augmented generation. xRAG reinterprets document embeddings in dense retrieval—traditionally used solely for retrieval—as features from the retrieval modality. By employing a modality fusion methodology, xRAG seamlessly integrates these embeddings into the language model representation space, effectively eliminating the need for their textual counterparts and achieving an extreme compression rate. In xRAG, the only trainable component is the modality bridge, while both the retriever and the language model remain frozen. This design choice allows for the reuse of offline-constructed document embeddings and preserves the plug-and-play nature of retrieval augmentation. Experimental results demonstrate that xRAG achieves an average improvement of over 10% across six knowledge-intensive tasks, adaptable to various language model backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts configuration. xRAG not only significantly outperforms previous context compression methods but also matches the performance of uncompressed models on several datasets, while reducing overall FLOPs by a factor of 3.53. Our work pioneers new directions in retrieval-augmented generation from the perspective of multimodality fusion, and we hope it lays the foundation for future efficient and scalable retrieval-augmented systems [^1] Code available at: https://github.com/Hannibal046/xRAG. .
Figure 1: xRAG enables efficient retrieval augmentation by adding one document token [X].
1 Introduction
Retrieval-Augmented Language Models (RALMs) [41, 21, 8, 13, 66] have demonstrated remarkable performance in various knowledge-intensive tasks. By retrieving domain-specific, long-tailed, and up-to-date knowledge from a non-parametric datastore, RALMs significantly extend the boundaries of parametric Large Language Models (LLMs). However, integrating entire documents into prompts can substantially increase inference costs and potentially exceed the context limits of LLMs [28, 75]. As shown in Figure 1, with the help of relevant document, LLM could generate the correct answer, albeit at the cost of processing a document that extends the original query by more than tenfold.
How can we mitigate the costs associated with extended context while maintaining the benefits of retrieval augmentation? Recent research interest has converged on a promising direction: Context Compression. This explores two main approaches: soft-prompting methods, such as Gist [56], AutoCompressor [14], and ICAE [19], which compress the context into dense memory slots, and hard-prompting methods, such as LLMLingua [28] and RECOMP [75], where compression is applied on the surface form. These approaches, however, either require significant memory for storing LLM activations (e.g., 1.05 MB per token as reported by [56]) or suffer from relatively low compression rates. More critically, these methods often fail to address a key aspect that sets RALMs apart from other NLP tasks: through large-scale contrastive learning with question-document pairs, modern dense retrieval systems already condense document information into a single high-dimensional embedding, and this embedding reveals (almost) as much information as the text [55, 43].
In this paper, we explore an interesting and novel perspective on retrieval augmentation and context compression—modality fusion. In multimodal literature where text-only language models can learn to "perceive" and "listen", a pretrained modality encoder, such as CLIP [60], is often utilized to extract modality features. These features are then integrated into language models through a modality fusion bridge [44, 50, 33]. Building on this conceptual similarity between retriever encoder and modality encoder, we propose xRAG, which reinterprets document embeddings in dense retrieval—traditionally used solely for retrieval—as features from the retrieval modality. By employing a modality fusion methodology, xRAG seamlessly integrates these embeddings into the language model representation space, eliminating the need for their textual counterparts and achieving an extreme compression rate. In xRAG, the only trainable component is the modality bridge, while both the retriever and the LLM remain frozen. This design choice allows for the reuse of offline-constructed document embeddings and preserves the plug-and-play nature of retrieval augmentation—two essential factors for a practical RAG system.
To verify the effectiveness and versality of our framework, we conducted comprehensive experiments with different LLM backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts model. Our results reveal that adding just one document token could lead to over a 10% improvement across six knowledge-intensive tasks, significantly surpassing previous compression methods. xRAG also delivers results comparable to uncompressed models on several datasets. This is remarkable considering that the only trainable component constitutes less than 0.1% of the LLM’s parameters. In terms of efficiency, xRAG reduces total FLOPs by a factor of 3.53 compared to the uncompressed RAG model. We further provide detailed analyses of xRAG, examining various training strategies, data blends, and component selections for the retrieval system. We hope this research lays the foundation for future efficient and scalable retrieval-augmented systems.
2 Related Work
Retrieval-augmented Generation
Equipping a parametric language model with a non-parametric datastore has proven effective for various NLP tasks, such as language modeling [36, 54, 81], open-domain question answering [24, 41, 66, 49], machine translation [35, 11] and others. Given the vast design space of this generation paradigm, numerous approaches with different focuses have been proposed. For instance, RETRO [8] and PlugLM [12] feature architectural innovations for better integration with the non-parametric datastore. REALM [21] introduces an end-to-end methodology for joint optimization of the language model and the retriever. REPLUG [66] and RA-DIT [48] enhance the retriever using LLM feedback to better align these two components. DSP [37] and InteR [17] explore sophisticated interactions between the retriever and the language model. Selfmem [13] employs a reward model to iteratively refine both retrieval and generation processes. Self-RAG [4] introduces a self-reflection mechanism that significantly enhances the quality and factuality of language models. For a more comprehensive review, refer to [18, 5, 3]. Our work, xRAG, distinguishes itself from previous studies by adopting a modality fusion approach to the problem of retrieval augmentation, resulting in a both effective and efficient RAG system.
Context Compression
Context compression, which aims to reduce the input length of LLMs while preserving the salient information of the original context, has gained significant attention recently. Gist [56] achieves up to a 26x compression rate of instruction by modifying the attention mask and caching the activation of soft gist tokens. ICAE [19] and AutoCompressor [14] compresses long contexts into short, compact memory slots that can be directly conditioned on by the LLM for various purposes. LLMLingua [28, 29] proposes a coarse-to-fine prompt compression method based on the perplexity score. These are general compressors, while others propose methods specifically designed for RAG, such as FilCo [72] and RECOMP [75]. A more detailed comparison between these compression methods in terms of memory efficiency, compression rate, and adaptability can be found in Appendix A.
Figure 2: Overview of xRAG (a) and RAG (b). For a given query, RAG typically concatenates the retrieved document with the query, significantly extending the context length. In contrast, xRAG addresses this issue through modality fusion by directly projecting the document embedding into the LLM’s representation space. This allows for efficient retrieval-augmentation with the addition of only one token.
3 Methods
Problem Formulation
[…]
3.1 xRAG Architecture
Traditional methods for document compression typically focus on surface form of the document [28, 72, 75]. In contrast xRAG tackle the problem from a modality fusion view. Concretely, we introduce a modality projector ${{\bf W}}$, which is trained to directly project the retrieval features E into the LLM representation space. Our proposed framework is visually contrasted with the traditional RAG system in Figure 2. In the standard RAG, the input to the LLM comprises the embeddings $\texttt{Emb}(\textrm{D}\oplus q)$ of length $|\textrm{D}|+|q|$, where Emb signifies the embedding layer of the LLM. Conversely, with xRAG, the modified input is represented as ${{\bf W}}(\textrm{E})\oplus\texttt{Emb}(q)$, which yields a substantially reduced length of $1+|q|$. In this framework, the challenges come from the modality fusion: How can a text-only language model understand features from retrieval modality? To achieve this, we explore a two-stage training strategy: Paraphrase Pretraining followed by Context-aware Instruction Tuning.
Figure 3: Two-stage training strategy of xRAG including (a) Paraphrase Pre-training on unlabeled corpus and (b) Context-aware Instruction Tuning optimized with labeled data and self-distillation.
3.2 Paraphrase Pretraining
[…]
$\mathcal{L}_{\text{nll}}=-\sum_{i=1}\text{log}\ p_{\phi}(d_{i}|{{\bf W}}(\textrm{E}),{{\bf X}}_{\texttt{instruct}},d_{<i})$ (1)
[…]
3.3 Context-aware Instruction Tuning
[…]
Optimization I: Language Modeling.
Aligned with established instruction tuning methodologies [71, 23, 48, 42], our objective is to finetune the model so that it generates the correct output when provided with a specific instruction, conditioned upon the given context information. Unlike traditional models that utilize the textual context ${{\bf X}}_{\texttt{context}}$, our method employs a dense feature $\textrm{E}_{\texttt{context}}$ to encapsulate the context information:
$\mathcal{L}_{\text{nll}}=-\sum_{i=1}\text{log}\ p_{\phi}({{\bf X}}_{\texttt{answer},i}|{{\bf W}}(\textrm{E}_{\texttt{context}}),{{\bf X}}_{\texttt{question}},{{\bf X}}_{\texttt{answer},<i})$ (2)
Optimization II: Self-Distillation.
[…]
$\mathcal{L}_{\text{kl}}=D_{\textrm{KL}}(p_{\phi}({{\bf X}}_{\texttt{answer}}|{{\bf X}}_{\texttt{context}},\cdot)\ ||\ p_{\phi}({{\bf X}}_{\texttt{answer}}|{{\bf W}}(\textrm{E}_{\texttt{context}}),\cdot))$ (3)
Here ${{\bf X}}_{\texttt{question}}$ is omitted for brevity and the final loss is the linear combination controlled by a hyperparameter: $\mathcal{L}_{nll}+\alpha\mathcal{L}_{kl}$.
3.4 Design Principle
In designing the projector $\mathbf{W}$, our primary objective is to maintain the simplicity of the framework. We therefore opted for a two-layer MLP while other more sophisticated module such as Q-Former [44] could also be considered. Notice that the projector is the only trainable component, accounting for only 0.46% of the total parameters in the Mistral-7b model and 0.07% in the Mixtral-8x7b model. Such a design choice departs from previous studies that necessitated full-parameter tuning to adapt LLMs for compressed contexts [56, 72, 14]. We believe this approach will likely be more accessible and practical because, fundamentally, the RAG itself functions as a plug-and-play module for LLMs, and so should its compressed version. This design also avoid the risk of compromising other core capabilities of LLM during full-parameter tuning, as observed in  [52, 53].
Moreover, in contrast to other compression methods that necessitate storing LLM activations for each compressed token [56, 19, 14]—an impractical strategy in the RAG setting, given the millions of documents involved—our method introduces no additional memory overhead. Instead, it leverages offline-constructed document embeddings, originally designed for retrieval. To summarize, xRAG not only simplifies the integration process but also avoids unnecessary computational or memory expenses.
4 Experimental Setup
4.1 Evaluation Dataset
We evaluated the performance of xRAG primarily on knowledge-intensive tasks which encompass a range of challenges: (1) three Open Domain Question Answering datasets that address questions on a wide array of topics: Natural Questions [40], TriviaQA [32], and Web Questions [7]. (2) one Multihop Question Answering dataset, HotpotQA [77], which necessitates multi-step reasoning to generate answers. (3) one Long-form Question Answering dataset, TruthfulQA [47], that requires the generation of long-form and truthful responses. (4) one fact-checking dataset, FactKG [38], which challenges the model to use complex reasoning to determine the factual accuracy of given claims.
In line with the KILT [59] and GenRead [79], we assessed three ODQA datasets and HotpotQA using the Exact Match (EM) metric, FactKG with Accuracy, and for the long-form QA, we used both the F1 score and Rouge-L (R-L) score. These tasks demand a broad spectrum of world knowledge and have been extensively explored in the retrieval-augmentation literature [41, 21, 34, 9, 24, 66].
4.2 Implementation Details
To demonstrate the versatility of our framework, we choose two backbones, differing in scale and architecture: Mistral-7b [26] and Mixtral-8x7b [27]. For the retrieval corpus, we utilized the Wikipedia dump from December 2021, which was pre-processed into passages following the methodology described in [25]. This resulted in approximately 37 million passages, each averaging 180 tokens in length. Our default retrieval model is the SFR [1], which, at the time of writing this paper, holds the leading position on the MTEB leaderboard [57]. We use top-1 ranked document for inclusion in our instruction-tuning dataset and for the evaluation of downstream tasks. More details are provided in Appendix D.
4.3 Baselines
In determining appropriate baselines for comparison, we adhered to a fundamental principle: the selected compression methods must support the general plug-and-play capability of retrieval augmentation. This entails that they should function effectively without the need for dataset-specific tuning [75, 72], or any alteration to the parameters of LLMs [72, 56]. Furthermore, given the extensive volume of the retrieval corpus, it is essential that these compression methods demonstrate memory efficiency, specifically by not requiring the storage of LLM activations for each individual token [56, 19, 14]. With these criteria in mind, our evaluation compares xRAG to the following baselines: (I) Primarily, we consider two variants of LLMs: one that operates without retrieval augmentation and another that includes it. These serve as the lower and upper performance bounds for our study of compression techniques, respectively. (II) Additionally, our comparisons extend to LLMLingua [28], a plug-and-play approach for context compression. (III) Taking inspiration from [56], we incorporate a method of discrete compression using TF-IDF. This approach yields compression rates comparable to those achieved by xRAG and serves as the lower bound for discrete compression.
NQ TriviaQA WebQA HotpotQA TrutufulQA FactKG Average # Doc Length
Task Type
Open-Domain QA
Multihop QA
Long-form QA
(F1/R-L)
Fact Checking
Mistral-7b
w/o retrieval 30.25 57.08 34.89 27.02 26.23 25.51 54.78 36.54 (0.0%) 0
w retrieval 42.71 65.88 37.84 38.79 26.50 25.92 67.76 43.63 (19.4%) 175.1
*with Compression
LLMLingua 30.64 57.94 32.63 29.91 25.70 25.10 64.17 38.01 (4.0%) 98.6
LLMLingua 28.81 57.09 32.33 29.13 26.10 25.39 63.57 37.48 (2.5%) 61.1
TF-IDF 30.25 58.49 35.43 26.62 26.33 25.83 59.56 37.49 (2.6%) 1
xRAG 39.10 65.77 39.40 34.05 28.10 27.71 63.08 42.46 (16.2%) 1
Mixtral-8x7b
w/o retrieval 41.99 71.10 40.31 32.87 25.60 24.90 62.64 42.76 (0.0%) 0
w retrieval 45.15 70.34 41.26 43.46 27.10 25.80 70.42 46.22 (8.0%) 175.1
*with Compression
LLMLingua 37.65 67.70 36.02 35.66 25.99 25.39 67.98 42.32 (-1.0%) 96.6
LLMLingua 37.81 67.81 35.78 35.27 25.68 25.00 68.03 44.17 (-1.3%) 61.1
TF-IDF 41.19 69.94 41.63 32.05 26.80 26.00 66.17 43.41 (1.4%) 1
xRAG 47.28 74.14 44.50 39.66 27.80 26.64 68.20 46.91 (9.7%) 1
Table 1: Experimental results on six downstream tasks. The best results are in bold and the second best are with underscore. Percentage in the brackets denotes the relative improvement over non-retrieval setting. LLMs are frozen during the experiments and retrieved documents are set the same for different compression methods. and denotes different compression ratio.
5 Experimental Results
5.1 Knowledge Intensive Tasks
In Table 1, we present our main results. Across both Mistral-7b and Mixtral-8x7b configurations, we observe a consistent and significant improvement when retrieval augmentation is applied (p-value < 0.05), although the gains are more modest for the larger model configurations. This trend aligns with observations reported by [66, 48]. Further analysis on the efficacy of various compression techniques reveals that xRAG outperforms other approaches by a large margin. Remarkably, xRAG not only reduces the token count drastically—from 175.1 to a single token—but also maintains robust performance levels. In some instances, xRAG’s performance is comparable to, or even exceeds, that of the uncompressed models. Specifically, in the Mistral-7b configuration, xRAG achieves nearly the same performance improvement as the uncompressed model (16.6% compared to 19.4%), and in the Mixtral-8x7b configuration, it surpasses the uncompressed model (9.7% compared to 8.0%). One possible reason lies in the vulnerability of current RAG system when the irrelevant or misleading documents are presented, a topic detailed discussed in §6.1. We also observe that xRAG performs well in tasks that require document understanding, such as TriviaQA. However, in tasks that demand reasoning over document, like HotpotQA and FactKG, xRAG lags behind by a considerable gap.
CUDA Time (ms) GFLOPs
RAG xRAG Improvement RAG xRAG Improvement
FactKG 431.5 215.6 x2.01 4683.8 1289.5 x3.63
NQ 918.7 611.3 x1.51 1338.6 384.0 x3.48
TriviaQA 807.1 512.1 x1.57 1667.2 492.3 x3.38
WebQA 872.6 577.3 x1.51 1405.1 386.8 x3.63
Average x1.64 x3.53
Table 2: Comparison of RAG and xRAG performance in CUDA Time and GFLOPS.
5.2 Computational Efficiency
In this section, we conduct a thorough assessment of our framework’s computational efficiency and memory management. To rigorously evaluate our model, we employed Torch Profiler [^3] https://pytorch.org/docs/stable/profiler.html#module-torch.profiler to measure the CUDA Time (milliseconds) and Giga FLOPs of both the RAG and xRAG models across four real-world datasets. In these evaluations, the Mistral-7b, operating in bfloat16 inference mode, served as the base LLM. CUDA Time and GFLOPs were calculated on an average per batch basis with a fixed batch size, and GFLOPs were normalized by the number of generated tokens. These experiments were performed on the same computational hardware, specifically an Nvidia A100 and an AMD EPYC 7V12 64-Core Processor. As depicted in Table 2, despite variations in prompt and generation lengths across the datasets, xRAG significantly outpaced the RAG model, achieving a x1.64 increase in CUDA Time efficiency and a x3.53 reduction in GFLOPs.
6 Analysis
6.1 Evaluation Beyond the Overall Score
Although retrieval augmentation generally boosts performance as shown by aggregate metrics, it may not uniformly benefit all instances. In certain cases, the retrieval system might provide irrelevant or even misleading information, leading to incorrect answers that were previously correct [78, 72, 5]. To enable a more fine-grained evaluation, we introduce two novel metrics: the Resilience Rate and the Boost Rate. The resilience rate quantifies the percentage of instances in which the system’s responses remain correct both before and after retrieval augmentation, highlighting the system’s stability and robustness. Conversely, the boost rate measures the percentage of instances that were initially answered incorrectly but were rectified following the introduction of a retrieved document, thereby assessing the efficacy of retrieval augmentation. An ideal RAG system should have both high resilience rate and boost rate.
In Figure 4, we display these metrics for the uncompressed RAG and two compression methods: LLMLingua and xRAG. Surprisingly, although retrieval augmentation generally enhances performance, the resilience rate for RAG averages only 75.2%, indicating that retrieval can adversely affect about one-quarter of previously correct answers. In contrast, xRAG demonstrates considerable robustness across all evaluated datasets. This robustness largely stems from xRAG’s ability to maintain an unbiased stance toward the internal knowledge representation of the LLM, especially when confronted with noisy retrieval content. Similar trends are noted in [48, 52], where search-augmented instruction learning is shown to bolster the robustness of language models. However, xRAG still lags behind RAG in boost rate, particularly in multi-hop reasoning tasks. It is crucial to note that a high resilience rate does not necessarily mean that the LLM disregards the provided information, which could potentially lead to a reduced boost rate. A comparative analysis with LLMLingua indicates that xRAG is not only more robust but also more effective.
Figure 4: Resilience rate and boost rate of three augmentation methods: LLMLinuga, xRAG and RAG over a Mixtral-8x7b baseline without retrieval augmentation.
6.2 What makes xRAG effective?
NQ TriviaQA WebQA HotpotQA Averaged
Performance Resilience Boost
Mistral-7b
xRAG 39.10 65.77 39.40 34.05 44.58 82.3% 22.2%
w/o finetune 30.14 59.48 35.19 26.70 37.87 66.6% 20.8%
w/o pretrain 31.25 59.07 41.19 24.32 38.95 79.8% 14.1%
w/o nll 35.46 65.27 39.57 31.80 43.02 83.7% 19.4%
w/o self-kd 34.99 64.33 39.22 27.45 41.49 76.2% 20.8%
w LoRA 35.71 60.14 40.45 22.91 39.80 76.0% 18.0%
Mixtral-8x7b
xRAG 47.48 74.14 44.50 39.66 51.45 84.9% 20.0%
w/o finetune 34.46 64.08 34.89 30.43 40.96 65.9% 17.8%
w/o pretrain 42.54 71.17 47.44 31.23 48.09 85.0% 14.2%
w/o nll 45.10 72.85 45.03 37.11 50.02 84.8% 18.9%
w/o self-kd 42.38 72.26 44.73 32.41 47.94 79.8% 18.9%
Table 3: Ablation on different training strategy for xRAG.
This section delves into a thorough evaluation of various elements that contribute to xRAG’s overall performance, focusing on its training strategy, the blend of datasets used and the effect of different embedding models. Due to the space limit, we present the last factor in Appendix E.
1. Training Strategy We carefully ablate four optimization choices: pretraining, instruction tuning, and two optimization objectives—language modeling (nll) and self-distillation (self-kd). We also train a Mistral-7b with LoRA on our instruction tuning dataset to rule out the possibility that our improvement simply comes from tuning on more data. The outcomes are presented in Table 3. Our analysis reveals that the interplay of different training strategies significantly contributes to the performance of our framework. In the case of Mistral-7b, pretraining and finetuning phases are of equal significance to the end results. However, for Mixtral-8x7b, the impact of pretraining is notably diminished, likely due to the larger model’s enhanced capability to incorporate multi-modality information. Furthermore, we find that during finetuning, self-distillation is more important than language modeling. The primary advantage of self-distillation lies in bolstering the resilience rate of the xRAG system. Optimization with nll loss tends to cause an overreliance on context information, rendering the system more vulnerable when the retriever fails to fetch relevant documents.
# Train NQ TriviaQA WebQA HotpotQA Average
Performance Resilience Boost
xRAG (Mistral-7b) 955k 39.10 65.77 39.40 34.05 44.58 82.3% 22.2%
w RC only 488k 36.98 65.77 41.39 32.82 44.24 81.9% 22.4%
w QA only 385k 36.45 65.57 41.14 31.80 43.74 80.5% 22.1%
w Summ only 81k 36.37 64.95 40.40 31.98 43.42 78.8% 22.8%
Table 4: Abaltion results on different data selection strategy.
Figure 5: Given the misleading document, RAG model tend to generate a wrong answer based on the document, while xRAG demonstrate its robustness by leveraging the internal knowledge of LLM.
II. Instruction-tuning Dataset Blend As discussed in §$\S$3.3, our instruction-tuning dataset primarily comprises three categories: reading comprehension, open-domain QA, and text summarization. To explore the effects of different data blends, we instruction-tune three xRAG model variants, each using data from these distinct categories. The results are shown in Table4. Our analysis reveals that among the dataset blends, reading comprehension data most significantly enhances the xRAG model’s performance, as evidenced by both high resilience and boost rates. Intriguingly, when tuned solely with summarization data, xRAG still manages to deliver strong performance on QA datasets it has never been exposed to. This finding underscores that the advantages of instruction tuning for xRAG are not rooted in task-specific knowledge. Instead, they derive from the model’s improved ability to utilize projected context information effectively.
6.3 Case Study
In Figure 5, we show one interesting case about the robustness of xRAG. When retrieval system provide misleading content, standard RAG would overly rely on the document and generate answer that are faithful to the document while not factually true. Our xRAG model opt to rely on the internal knowledge of LLM and being robust to the misleading content. In Appendix H, we include more cases about xRAG including several error analysis.
7 Conclusion
In this work, we present xRAG, an innovative context compression method tailored for retrieval-augmented generation. For knowledge-intensive tasks, xRAG can be significantly faster than RAG while maintaining comparable performance. We are excited about the future of this modality-based retrieval-augmented system and plan to further improve its performance in the areas of reasoning over embedding, handling multiple documents, and combining with multi-vector retrieval.
Appendix A Comparison between different Context Compression Models
In Table 5, we present a detailed comparison of various context compression models, emphasizing their real-world applicability. This comparison focuses on two key aspects: (1) Plug-and-play capability, which assesses whether dataset-specific tuning is necessary for new, unseen data; (2) Memory efficiency, which evaluates if additional memory space is required to store the compressed information, such as high-dimensional vectors typically used in soft prompting methods.
Model
Specifically
designed for RAG
Maximun
Compression Rate
Approach Plug-and-Play Memory Efficient
AutoCompressor [14] x15 Soft Prompting
Gist [56] x26 Soft Prompting
ICAE [19] x8 Soft Prompting
LLMLingua [28] x20 Prompt Editing
Selective Context [45] x5 Prompt Editing
Token Elimination [80] x10 Attention Filtering
FilCo [70] x2 Prompt Editing
RECOMP [75] x16.6 Prompt Editing
xRAG x178 Modality Fusion
Table 5: Comparison between different compression methods from their setting to design principle.
Appendix B Instructions for Paraphrase Pretraining
The list of instructions for Paraphrase Pretraining is shown in Table 6. They present the same meaning with natural language variance.
"Background: [X] means the same as [D] " "Background: [X] Can you put the above sentences in your own terms? [D] " "[X] Please provide a reinterpretation of the preceding background text. [D] " "These two expressions are equivalent in essence:(1) [X] (2) [D] " "Background: [X] is a paraphrase of what? [D] " "[X] Could you give me a different version of the background sentences above? [D] " "In other words, background: [X] is just another way of saying: [D] " "You’re getting across the same point whether you say background: [X] or [D] " "[X] After uppacking the ideas in the background information above, we got: [D] " "[X] Please offer a restatement of the background sentences I’ve just read. [D] " "Background: [X] , which also means: [D] " "Strip away the mystery, and you’ll find [X] is simply another rendition of: [D] " "The essence of background: [X] is captured again in the following statement: [D] "
Table 6: Instructions used for Paraphrase Pretraining where [X] and [D] are placeholders for projected retrieval feature W(E) and document D.
Appendix C Details for Instruction Tuning Dataset
After collecting the raw data from different categories, we use templates [^4] https://github.com/google-research/FLAN/blob/main/flan/templates.py from FLAN [73] to construct instruction tuning dataset. In Table 7 we list an overview and in Table 8 we list the detailed information for each subtask of our dataset. For QA tasks that lack an explicit context, we perform a retrieval operation within the corpus $\mathbb{D}$ to identify the most relevant document to serve as context. This approach is akin to the retrieval-augmented instruction tuning depicted in [52, 48].
Task Type # Involved datasets # Train # Prompt # Label
Reading Comprehension 7 488,344 447.62 30.34
Summarization 3 81,821 483.49 53.29
Open Domain QA 7 385,173 203.55 20.09
Table 7: Overall statistics of Instruction Tuning dataset.
Task Type Dataset # Train # Prompt Len # Label Len
Reading Comprehension CoQA [62] 7101 617.98 77.75
DROP [16] 76098 356.06 3.86
NarrativeQA [39] 32747 702.39 7.86
PubMedQA [31] 1000 397.91 65.4
QuAIL [63] 10246 512.9 2.0
SQuAD v2 [61] 130319 214.54 6.87
PwC [19] 241564 571.35 53.07
Open Domain QA NQ [40] 87925 203.62 5.976
TriviaQA [32] 78785 216.1 6.49
CommonsenseQA [68] 9741 223.64 2.0
WikiQA [76] 1040 192.89 40.79
YahooQA [^5] https://huggingface.co/datasets/yahoo_answers_qa 87358 196.56 56.7
FreebaseQA [30] 20353 218.49 4.87
MSMarco [6] 99994 194.82 15.91
Summarization CNN/DM [65] 100000 616.99 63.37
SamSum [20] 14731 187.87 29.12
DialogSum [10] 12460 247 37.61
Table 8: Detailed data statistics for our Context-aware Instruction Tuning Dataset.
Appendix D Implementation Details
For the language models we use, Mixtral-8x7b is approximately 6.5 times larger in scale compared to Mistral-7b and features a divergent architectural approach—specifically, a dense versus mixture-of-experts design. For our assessments, we employed the instruction-tuned variants of these models.
Owing to efficiency constraints, we opted not to perform on-the-fly retrieval. Instead, we pre-constructed a retrieval index using the efficient and robust multi-vector retriever, ColBERT-v2 [64], from which we retrieved the top-1 ranked document for inclusion in our instruction-tuning dataset and for the evaluation of downstream tasks. Subsequently, we re-encoded these documents using the embedding model of interest (e.g., SFR). This strategy allows us to iterate data-centric experiments quickly. All experiments are conducted on the a setup of 8xNvidia A100 GPUs.
In Table 9 and Table 10, we list the hyperparameters for Paraphrase Pretraining and Context-aware Instruction Tuning.
Hyperparameter Assignment
optimizer AdamW
learning rate 6e-3
lr scheduler type linear
warmup ratio 0.03
weight dacay 0.0
epochs 1
flash attention True
batch size 12
gradient accumulation steps 4
num GPUs 8
max sequence length 336
max train samples 2,000,000
Table 9: Hyperparameters for Paraphrase Pretraining.
Hyperparameter Assignment
optimizer AdamW
learning rate 2e-5
lr scheduler type linear
warmup ratio 0.03
weight dacay 0.0
epochs 1
KL $\alpha$ 2.0
KL temperature 1.0
flash attention True
batch size 4
gradient accumulation steps 2
num GPUs 8
max sequence length 1024
max train samples 955,338
Table 10: Hyperparameters for Context-aware Instruction Tuning.
Appendix E About different Embedding Models
In our primary experiments, we use the SFR model as our default sentence embedding model. This section delves into the effects of different embedding models. We examine four universal text embedding models: E5-Mistral and E5-Large [69] alongside BGE-Large and BGE-Base [74]. Additionally, we assess two retrieval-specific models: Dragon [46] and DPR [34]. The configurations of the different retrievers and their MTEB scores [^6] https://huggingface.co/spaces/mteb/leaderboard—a metric indicating their general sentence representation capability—are listed in Table E. To isolate the impact of potentially different retrieved documents, we ensure that all models utilize the same top-1 document. The performance is averaged over four question answering datasets. A general pattern is that embedding models with stronger sentence representation capabilities tend to further enhance the downstream performance. Remarkably, the Dragon model, despite being a BERT-base-sized retrieval-specific model, outperforms general text embedding models that are twice its size (BGE-Large).
Model Model Size Embedding Dim Universal Embedding MTEB Score Average
Performance Resilience Boost
Mistral-7b
w/o retrieval 37.2 - -
w retrieval 46.3 74.1% 29.5%
[…]
w E5-Mistral 7B 4096 66.63 44.0 84.0% 20.6%
w E5-Large 335M 1024 62.25 42.1 80.2% 19.6%
w BGE-Large 335M 1024 64.23 41.6 78.2% 19.8%
w BGE-Base 109M 768 63.55 41.2 78.7% 18.9%
w Dragon 109M 768 - 42.1 84.2% 16.9%
w DPR 109M 768 - 40.5 77.4% 18.2%
Table 11: Ablation on different sentence embedding models.
Appendix F Analysis on Mistral-7b
In Figure 6, we list the Resilience rate and Boost Rate on Mistral-7b model, which exhibit same pattern with Mixtral-8x7b model.
Figure 6: Robustness and effectiveness analysis on 4 QA datasets with Mistral-7b model.
Appendix G Limitations
We discuss the limitations of our framework as follows:
  • In this work, we only consider the most commonly used retrieval system—single dense vector retrieval, while sparse retrieval methods such as BM25 or multi-vector retrieval methods like ColBERT are not included. We believe that combining these methods would be a promising direction for xRAG, as sparse vectors could complement dense vectors, and multi-vector retrieval would provide xRAG with more flexibility by not condensing all information into one token.
  • Currently, xRAG delivers decent performance when a relevant document is fetched; however, it lags behind RAG by a considerable margin in tasks that require reasoning (such as HotpotQA and FactKG). One possible reason is that during the training phase of xRAG, reasoning-relevant data is not provided. How to make xRAG a better reasoner remains our future work.
  • We only consider the Top-1 retrieval setting, while ensembling multiple relevant documents has been shown to be effective for RAG systems due to the complementary information contained in Top-K documents. We believe there is potential advantage for xRAG to scale to multi-document settings, as the input length of xRAG for multi-documents scales by a factor of 1, while for RAG, it scales by the document length factor.
Appendix H More interesting cases
  • In Figure 7, we report a failure case of xRAG. In this case, retrieval alone is not enough to derive the final answer and the LLM is required to perform reasoning over retrieved document (the listed universities are all located in Switzerland).
  • An interesting example is shown in Figure 8, when the retrieved document is a list of a characters in the book Discworld, the RAG model would respond with a fictional character, while xRAG generate the right answer by focusing on the relevant part of the document.
  • In Figure 9, even when the retrieved document is relevant, RAG would still hallucinate while xRAG could generate the right answer based on the document.
  • In Figure 10, the retriever mistakenly fetch the wrong document (Phantom of the Opera of interest is a music rather than a file) and RAG would be misled while xRAG remain robust to generate the correct answer.
Figure 7: Failure case of xRAG when reasoning is required to derive the final answer.
Figure 8: xRAG correctly locates the relevant part in a long document by selecting Diogenes of Sinope as the answer rather than Didactylos, a fictional character in the book Discworld.
Figure 9: xRAG correctly locates the relevant part in a long document while RAG would still hallucinate the wrong answer.
Figure 10: xRAG correctly locates the relevant part in a long document while RAG would still hallucinate the wrong answer.

### v2

@@
 Abstract
-This paper introduces xRAG, an innovative context compression method tailored for retrieval-augmented generation.
-xRAG reinterprets document embeddings in dense retrieval—traditionally used solely for retrieval—as features from the retrieval modality.
-By employing a modality fusion methodology, xRAG seamlessly integrates these embeddings into the language model representation space, effectively eliminating the need for their textual counterparts and achieving an extreme compression rate.
-In xRAG, the only trainable component is the modality bridge, while both the retriever and the language model remain frozen.
+This paper introduces xRAG, a novel context compression method designed specifically for retrieval-augmented generation.
+xRAG redefines the use of document embeddings in dense retrieval—traditionally limited to retrieval purposes—by integrating them as features from the retrieval modality.
+Through a modality fusion approach, xRAG effectively merges these embeddings into the language model’s representation space, eliminating the need for their textual counterparts and achieving an extreme compression rate.
+In xRAG, the modality bridge is the only trainable component, while the retriever and language model remain frozen.
 This design choice allows for the reuse of offline-constructed document embeddings and preserves the plug-and-play nature of retrieval augmentation.
-Experimental results demonstrate that xRAG achieves an average improvement of over 10% across six knowledge-intensive tasks, adaptable to various language model backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts configuration.
-xRAG not only significantly outperforms previous context compression methods but also matches the performance of uncompressed models on several datasets, while reducing overall FLOPs by a factor of 3.53.
-Our work pioneers new directions in retrieval-augmented generation from the perspective of multimodality fusion, and we hope it lays the foundation for future efficient and scalable retrieval-augmented systems [^1] Code available at: https://github.com/Hannibal046/xRAG.
-.
+Experimental results demonstrate that xRAG achieves an average improvement of over 10% across six knowledge-intensive tasks, compatible with various language model backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts configuration.
+xRAG not only significantly outperforms previous context compression methods but also matches the performance of uncompressed models on several benchmarks, while reducing overall FLOPs by a factor of 3.53.
+This work pioneers new avenues in retrieval-augmented generation through multimodal fusion, potentially setting a groundwork for future developments in efficient and scalable retrieval systems.
 Figure 1: xRAG enables efficient retrieval augmentation by adding one document token [X].
 1 Introduction
-Retrieval-Augmented Language Models (RALMs) [41, 21, 8, 13, 66] have demonstrated remarkable performance in various knowledge-intensive tasks.
+Retrieval-Augmented Language Models (RALMs) [41, 21, 8, 13, 70] have shown exceptional performance in a variety of knowledge-intensive tasks.
 By retrieving domain-specific, long-tailed, and up-to-date knowledge from a non-parametric datastore, RALMs significantly extend the boundaries of parametric Large Language Models (LLMs).
-However, integrating entire documents into prompts can substantially increase inference costs and potentially exceed the context limits of LLMs [28, 75].
-As shown in Figure 1, with the help of relevant document, LLM could generate the correct answer, albeit at the cost of processing a document that extends the original query by more than tenfold.
-How can we mitigate the costs associated with extended context while maintaining the benefits of retrieval augmentation?
+However, the integration of entire documents into prompts can significantly increase inference costs and may surpass the context limit of LLMs [28, 79].
+As illustrated in Figure 1, while the inclusion of a relevant document enables the LLM to generate accurate responses, it does so at the expense of processing documents that expand the original query by more than tenfold.
+How might we mitigate the costs associated with extended context while maintaining the benefits of retrieval augmentation?
 Recent research interest has converged on a promising direction: Context Compression.
-This explores two main approaches: soft-prompting methods, such as Gist [56], AutoCompressor [14], and ICAE [19], which compress the context into dense memory slots, and hard-prompting methods, such as LLMLingua [28] and RECOMP [75], where compression is applied on the surface form.
+This concept is pursued through two primary strategies: soft-prompting methods, such as Gist [58], AutoCompressor [14], and ICAE [19], which compress the context into dense memory slots, and hard-prompting methods, such as LLMLingua [28] and RECOMP [79], where compression is applied on the surface form.
 These approaches, however, either require significant memory for storing LLM activations (e.g., 1.05 MB per token as reported by [58]) or suffer from relatively low compression rates.
-More critically, these methods often fail to address a key aspect that sets RALMs apart from other NLP tasks: through large-scale contrastive learning with question-document pairs, modern dense retrieval systems already condense document information into a single high-dimensional embedding, and this embedding reveals (almost) as much information as the text [55, 43].
-In this paper, we explore an interesting and novel perspective on retrieval augmentation and context compression—modality fusion.
-In multimodal literature where text-only language models can learn to "perceive" and "listen", a pretrained modality encoder, such as CLIP [60], is often utilized to extract modality features.
-These features are then integrated into language models through a modality fusion bridge [44, 50, 33].
-Building on this conceptual similarity between retriever encoder and modality encoder, we propose xRAG, which reinterprets document embeddings in dense retrieval—traditionally used solely for retrieval—as features from the retrieval modality.
-By employing a modality fusion methodology, xRAG seamlessly integrates these embeddings into the language model representation space, eliminating the need for their textual counterparts and achieving an extreme compression rate.
-In xRAG, the only trainable component is the modality bridge, while both the retriever and the LLM remain frozen.
-This design choice allows for the reuse of offline-constructed document embeddings and preserves the plug-and-play nature of retrieval augmentation—two essential factors for a practical RAG system.
-To verify the effectiveness and versality of our framework, we conducted comprehensive experiments with different LLM backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts model.
+More critically, these methods overlook a crucial characteristic of RALMs: through large-scale contrastive learning with question-document pairs, modern dense retrieval systems already distill document content into a single high-dimensional embedding, and this embedding reveals (almost) as much information as the text [57, 43].
+In this paper, we pioneer an innovative approach to retrieval augmentation and context compression through the lens of modality fusion.
+Drawing from multimodal research, where text-only language models are taught to "perceive" and "listen," a pretrained modality encoder like CLIP [64] is typically used to extract modality features.
+These features are then integrated into language models using a modality fusion bridge [44, 51, 33].
+Building on the conceptual overlap between the retriever encoder and modality encoder, we introduce xRAG.
+This model redefines document embeddings from dense retrieval—traditionally solely for retrieval purposes—as retrieval modality features.
+xRAG employs a modality fusion methodology to seamlessly integrate these embeddings into the language model’s representation space, thus obviating the need for textual counterparts and achieving significant context compression.
+In xRAG, the modality bridge is the only trainable component, while both the retriever and the LLM are kept frozen.
+This design decision facilitates the reuse of pre-constructed document embeddings and maintains the plug-and-play nature of retrieval augmentation—two essential factors for a functional RAG system.
+To verify the effectiveness and versatility of our framework, we conducted comprehensive experiments with different LLM backbones, ranging from a dense 7B model to an 8x7B Mixture of Experts model.
 Our results reveal that adding just one document token could lead to over a 10% improvement across six knowledge-intensive tasks, significantly surpassing previous compression methods.
-xRAG also delivers results comparable to uncompressed models on several datasets.
+xRAG also delivers results comparable to uncompressed models on several benchmarks.
 This is remarkable considering that the only trainable component constitutes less than 0.1% of the LLM’s parameters.
@@
 We further provide detailed analyses of xRAG, examining various training strategies, data blends, and component selections for the retrieval system.
-We hope this research lays the foundation for future efficient and scalable retrieval-augmented systems.
+We believe this research sets a strong foundation for the development of future efficient and scalable retrieval-augmented systems.
 2 Related Work
 Retrieval-augmented Generation
-Equipping a parametric language model with a non-parametric datastore has proven effective for various NLP tasks, such as language modeling [36, 54, 81], open-domain question answering [24, 41, 66, 49], machine translation [35, 11] and others.
+Equipping a parametric language model with a non-parametric datastore has proven effective for a range of NLP tasks, including language modeling [36, 56, 87], open-domain question answering [24, 41, 70, 86], domain adaptation [6] and machine translation [35, 11], among others.
 Given the vast design space of this generation paradigm, numerous approaches with different focuses have been proposed.
-For instance, RETRO [8] and PlugLM [12] feature architectural innovations for better integration with the non-parametric datastore.
-REALM [21] introduces an end-to-end methodology for joint optimization of the language model and the retriever.
-REPLUG [66] and RA-DIT [48] enhance the retriever using LLM feedback to better align these two components.
-DSP [37] and InteR [17] explore sophisticated interactions between the retriever and the language model.
-Selfmem [13] employs a reward model to iteratively refine both retrieval and generation processes.
-Self-RAG [4] introduces a self-reflection mechanism that significantly enhances the quality and factuality of language models.
-For a more comprehensive review, refer to [18, 5, 3].
-Our work, xRAG, distinguishes itself from previous studies by adopting a modality fusion approach to the problem of retrieval augmentation, resulting in a both effective and efficient RAG system.
+For instance, RETRO [8] and PlugLM [12] introduce architectural innovations for enhanced integration with the non-parametric datastore.
+REALM [21] pioneers an end-to-end approach for simultaneous optimization of the language model and retriever.
+REPLUG [70] and RA-DIT [50] improve retriever alignment using feedback from LLMs.
+DSP [37] and InteR [17] investigate complex interactions between the retriever and the language model.
+Selfmem [13] utilizes a reward model to refine retrieval and generation iteratively.
+Self-RAG [3] incorporates a self-reflection mechanism to enhance the quality and factuality of language model outputs.
+For a detailed overview, see [18, 4, 2].
+Our contribution, xRAG, stands out by implementing a modality fusion approach to retrieval augmentation, creating an effective and efficient RAG system.
 Context Compression
-Context compression, which aims to reduce the input length of LLMs while preserving the salient information of the original context, has gained significant attention recently.
-Gist [56] achieves up to a 26x compression rate of instruction by modifying the attention mask and caching the activation of soft gist tokens.
-ICAE [19] and AutoCompressor [14] compresses long contexts into short, compact memory slots that can be directly conditioned on by the LLM for various purposes.
-LLMLingua [28, 29] proposes a coarse-to-fine prompt compression method based on the perplexity score.
-These are general compressors, while others propose methods specifically designed for RAG, such as FilCo [72] and RECOMP [75].
-A more detailed comparison between these compression methods in terms of memory efficiency, compression rate, and adaptability can be found in Appendix A.
+Context compression, aimed at reducing the input length for LLMs while retaining essential information, has recently attracted substantial interest [46].
+Gist [58] achieves a compression rate of up to 26x by modifying the attention mask and caching soft gist token activations.
+ICAE [19]AutoCompressor [14], and 500xCompressor[47] condense lengthy contexts into succinct, compact memory slots, which are directly utilizable by LLMs for diverse functions.
+LLMLingua [28, 29, 62] and CompAct[82] introduces a coarse-to-fine prompt compression technique based on perplexity scores and distilled token-level score.
+While these methods are generally applicable, others are tailored specifically for RAG systems, such as FilCo [76] and RECOMP [79].
+A concurrent work directly employs passage embeddings for efficient listwise reranking [52].
+For an in-depth comparison of these compression methods regarding memory efficiency, compression rates, and adaptability, refer to Appendix A.
 Figure 2: Overview of xRAG (a) and RAG (b).
@@
 More details are provided in Appendix D.
+Code is available at: https://github.com/Hannibal046/xRAG.
 4.3 Baselines
@@
 This approach yields compression rates comparable to those achieved by xRAG and serves as the lower bound for discrete compression.
+Table 1: Experimental results on six downstream tasks.
+The best results are in bold and the second best are with underscore.
+Percentage in the brackets denotes the relative improvement over non-retrieval setting.
+LLMs are frozen during the experiments and retrieved documents are set the same for different compression methods.
+and denotes different compression ratio.
 NQ TriviaQA WebQA HotpotQA TrutufulQA FactKG Average # Doc Length
@@
 xRAG 47.28 74.14 44.50 39.66 27.80 26.64 68.20 46.91 (9.7%) 1
-Table 1: Experimental results on six downstream tasks.
-The best results are in bold and the second best are with underscore.
-Percentage in the brackets denotes the relative improvement over non-retrieval setting.
-LLMs are frozen during the experiments and retrieved documents are set the same for different compression methods.
-and denotes different compression ratio.
 5 Experimental Results
@@
 However, in tasks that demand reasoning over document, like HotpotQA and FactKG, xRAG lags behind by a considerable gap.
+Table 2: Comparison of RAG and xRAG performance in CUDA Time and GFLOPS.
 CUDA Time (ms) GFLOPs
@@
 Average x1.64 x3.53
-Table 2: Comparison of RAG and xRAG performance in CUDA Time and GFLOPS.
 5.2 Computational Efficiency
@@
 6.2 What makes xRAG effective?
+Table 3: Ablation on different training strategy for xRAG.
 NQ TriviaQA WebQA HotpotQA Averaged
@@
 w/o self-kd 42.38 72.26 44.73 32.41 47.94 79.8% 18.9%
-Table 3: Ablation on different training strategy for xRAG.
 This section delves into a thorough evaluation of various elements that contribute to xRAG’s overall performance, focusing on its training strategy, the blend of datasets used and the effect of different embedding models.
 Due to the space limit, we present the last factor in Appendix E.
+Figure 5: Given the misleading document, RAG model tend to generate a wrong answer based on the document, while xRAG demonstrate its robustness by leveraging the internal knowledge of LLM.
 1.
@@
 Optimization with nll loss tends to cause an overreliance on context information, rendering the system more vulnerable when the retriever fails to fetch relevant documents.
+Table 4: Abaltion results on different data selection strategy.
 # Train NQ TriviaQA WebQA HotpotQA Average
@@
 w Summ only 81k 36.37 64.95 40.40 31.98 43.42 78.8% 22.8%
-Table 4: Abaltion results on different data selection strategy.
-Figure 5: Given the misleading document, RAG model tend to generate a wrong answer based on the document, while xRAG demonstrate its robustness by leveraging the internal knowledge of LLM.
 II.
@@
 We are excited about the future of this modality-based retrieval-augmented system and plan to further improve its performance in the areas of reasoning over embedding, handling multiple documents, and combining with multi-vector retrieval.
+8 Acknowledgement
+We would like to express our sincere gratitude to the anonymous reviewers for their thorough review, insightful comments, and constructive suggestions, which have significantly improved the quality of this manuscript.
+This work paper is supported (in part) by the State Key Laboratory of General Artificial Intelligence.
 Appendix A Comparison between different Context Compression Models
@@
 Figure 10: xRAG correctly locates the relevant part in a long document while RAG would still hallucinate the wrong answer.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Abstract and Section 1
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 5.1 and Appendix G
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: the paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: in Appendix D we provide full details about our training process
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: all the models and data we use are publicly available and we carefully cite each paper
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Section 4.1
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: in Section 5.1
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: in Appendix D
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: reviewed and confirmed
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: in section 5.2
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: all the data and model we use is publicly available
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer:[N/A]
+  59. Justification: in Section C
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: n/a
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: n/a
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: n/a
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.